# Automated Explanation Selection for Scientific Discovery

### v2

@@ 3 Desirable Properties of Explanations
Explanation selection is also closely related to the notion of abduction, which was introduced by Pierce as a third mode of reasoning in addition to deduction and induction. Deduction is the process of deriving a result from a known fact and a rule. Induction, on the other hand, is the process of generalizing a rule from a set of observed facts and results. Abduction, on the other hand, is the process of proposing a fact to explain a known result using a known rule [33].
There is a large body of work on abductive reasoning in the context of the philosophy of logic and scientific discovery [33]. Pierce developed a theory of science in which abduction is the step in which a hypothesis is generated, i.e., a fact that explains an observation is suggested by scientific intuition, then deduction is used to evaluate the plausibility of the hypothesis, and then induction is used to provide further evidence for the hypothesis through experimentation.
In AI research, interpretations that capture abduction as a theory of “Inference to the Best Explanation” (IBE) are more common [28]. In the epistemological framework of Hintikka, the central concept is an interrogative model of inquiry, in which deduction plays a significant role in the validation of explanations [17].
@@ 5 Conclusion
We have proposed a cycle of scientific discovery that combines machine learning with automated reasoning for the generation and the selection of explanations. Our findings indicate that automated reasoning methods are well-suited for the generation of explanations for machine learning models, and that they can be used to guide the selection of explanations. We have also introduced a taxonomy of explanation selection problems that draws on insights from sociology and cognitive science. These selection criteria subsume existing notions and extend them with new properties. We discussed how the evaluation of explanations is greatly simplified through the guarantees provided by automated reasoning methods. We conclude that the application of automated reasoning methods to the selection of explanations for machine learning models represents a promising avenue for future research.

### v4

@@ 3 Desirable Properties of Explanations
Explanation selection is also closely related to the notion of abduction, which was introduced by Peirce as a third mode of reasoning in addition to deduction and induction. Deduction is the process of deriving a result from a known fact and a rule. Induction, on the other hand, is the process of generalizing a rule from a set of observed facts and results. Abduction, on the other hand, is the process of proposing a fact to explain a known result using a known rule [34].
There is a large body of work on abductive reasoning in the context of the philosophy of logic and scientific discovery [34]. Peirce developed a theory of science in which abduction is the step in which a hypothesis is generated, i.e., a fact that explains an observation is suggested by scientific intuition, then deduction is used to evaluate the plausibility of the hypothesis, and then induction is used to provide further evidence for the hypothesis through experimentation.
In AI research, interpretations that capture abduction as a form of “Inference to the Best Explanation” (IBE) are common [29]. In contrast, Hintikka’s notion of abduction is embedded in an interrogative model of inquiry [18] and understood as the introduction of new explanatory hypotheses, as opposed to IBE’s reasoning with a fixed hypothesis space [17]. This view is closer to Peirce’s notion of abduction, emphasizing hypothesis formation over selection among given alternatives; whereas IBE reduces to deduction over a fixed set of hypotheses with additional preference criteria.
@@ 5 Conclusion
We proposed a cycle of scientific discovery that combines machine learning with automated reasoning for generating and selecting explanations. Our findings indicate that automated reasoning methods are well-suited for generating explanations for machine learning models and can guide their selection. We also introduced a taxonomy of explanation selection problems, drawing on insights from sociology and cognitive science, which subsumes existing notions and extends them with new properties. We showed that evaluation is greatly simplified through the guarantees provided by automated reasoning methods. Overall, applying automated reasoning to explanation selection for machine learning models represents a promising direction for future research.